OpenStudy (anonymous):

Find the equation of the tangent line to F(x)=(x-8)^4 at the point where x = 2

myininaya (myininaya):

So you need to use chain rule to find F'

myininaya (myininaya):

Have you done that?

OpenStudy (anonymous):

yea i got 4(x-8)^3

myininaya (myininaya):

ok great now find F'(2) to find the slope at (2,F(2))

OpenStudy (anonymous):

not sure how to set that equation up

myininaya (myininaya):

You told me F'(x) which is 4(x-8)^3 So we have F'(x)=4(x-8)^3 Since you want to find F'(2) you just replace x with 2 in F'(x)=4(x-8)^3 like so F'(2)=4(2-8)^3=....

myininaya (myininaya):

\[F'(2)=4(2-8)^3=4(-6)^3=-4(6)^3\] -4(6)^3 is the slope at (2,F(2)) And by the way F(2)=(2-8)^4=(-6)^4=6^4

myininaya (myininaya):

\[y=mx+b \] is the form in which all lines can be written for this line we have m=-4(6)^3 and the point in which the line goes through which is (2,6^4)

myininaya (myininaya):

\[6^4=-4(6)^3 \cdot 2+b\] I replaced (x,y) with (2,6^4) and m with -4(6)^3 to find the y-intercept of the tangent line b

myininaya (myininaya):

\[y=-4(6)^3x+(6^4+4(6)^3 \cdot 2)\]